# Beyond Check-in Counts: Redefining Popularity for POI Recommendation with Users and Recency

### v1

@@ 5.1. Dataset
Note: The GETNext model’s output is shown in the first row. The performance metrics for various combinations of $\alpha$ and $\beta$ parameters are shown in the subsequent rows. The exact values of $\alpha$ and $\beta$ for each matching row are shown in the first two columns, respectively.
5.2. Evaluation Metrics
[…]
$\text{MRR}=\frac{1}{m}\sum_{i=1}^{m}\frac{1}{\text{rank}}$
5.3. Results
We have conducted a number of well designed experiments that have contributed significantly to our understanding of the complex dynamics of our model and marked our journey through the dataset. Despite the trials, the fundamental pillars of our study, denoted by the mysterious variables $\alpha$ and $\beta$, as stated in equation 1, held firm. Our technique stayed largely unchanged across a resolute span of 20 epochs, which allowed our comparison study to remain coherent.
As we carefully examined the temporal structure of our dataset, an important distinction became apparent. Check-ins sought solace in two discrete temporal domains—the recent and the past—amid the complex pattern of temporal dynamics. Those who were recorded in the history of the previous three months were labeled as recent. On the other hand, those who came before this point in time were relegated to the past.
6. Conclusion
We set out to redefine the concept of popularity assigned to a point-of-interest (POI) within the constraints of this research in order to further the area of POI recommendation systems. Our work is completed by integrating this new idea into GETNext, a unique graph-based learning model designed for POI recommendation tasks on a global graph structure. Elaborating on the complex relationship between the model’s effectiveness and the subtle interactions between two important parameters, $\alpha$ and $\beta$, as shown by equation 1, we approached this task with great care. We thoroughly examined the model’s performance dynamics over a range of $\alpha$ and $\beta$ values using a stringent set of experiments on an extensive real-world dataset, revealing a wealth of information about the complex behavior of the recommendation system. But our investigation doesn’t finish here; rather, it’s just the beginning of a longer journey to learn more about the time aspect of POI recommendation. We will next attempt to understand the complex time dynamics at work, determine the time influence on other variables, and clarify how the more general travel patterns change with the seasons. A greater comprehension of the fundamental processes guiding POI recommendation systems is anticipated to be revealed by these upcoming studies, opening the door to more reliable and contextually sensitive recommendation frameworks.

### v2

@@ 5.3 Results
Note: The GETNext model’s output is shown first as baseline. The performance metrics for various combinations of $\alpha$ and $\beta$ parameters are shown in the subsequent rows. The exact values of $\alpha$ and $\beta$ for each matching row are shown in the first two columns, respectively.
6 Conclusion
Our aim was to further the field of POI recommendation systems by redefining the definition of popularity ascribed to a point-of-interest (POI) within the limitations of this study. Our interpretation of POI popularity guarantees that recommended POIs represent the tastes of a larger, more diversified user base, rather than being biased by a small number of very active users, by taking into account both the check-in count and the number of unique users. Also, recommendations are kept up to date with the latest user behavior thanks to the integration of recent trends.
POI recommendation systems will be significantly impacted by this strategy. By avoiding biases brought about by excessively frequent check-ins from a small number of people, it improves fairness. To further represent the interests of the larger community, the model encourages recommendations that are more socially inclusive. We implement this new idea into a graph-based learning model called GETNext, which enables the scalable implementation of POI suggestions on a global graph structure. Strong and contextually relevant recommendations are ensured by the interplay between parameters $\alpha$ and $\beta$, as described in equation 1. This further refines the balance between user diversity and current trends.
Our investigation doesn’t finish here; rather, it’s just the beginning of a longer journey to learn more about the time aspect of POI recommendation. We will next attempt to understand the complex time dynamics at work, determine the time influence on other variables, and clarify how the more general travel patterns change with the seasons. A greater comprehension of the fundamental processes guiding POI recommendation systems is anticipated to be revealed by these upcoming studies, opening the door to more reliable and contextually sensitive recommendation frameworks.